Status Report

ISS On-Orbit Status 2 Aug 2002

By SpaceRef Editor
August 2, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted

The crew was thanked for squeezing in some last minute robotic
activities before dinner last evening.  This was instrumental in
confirming the source of one of several error indications.

During yesterdayâs robotics operations, not all SSRMS (space station
remote manipulator system) and MBS (mobile base system) checkout
objectives were completed.  The remaining OCRs (on-orbit
checkout requirements) will be added to the next MSS (mobile service
system) ops day.  The SSRMS will remain in the POA viewing
position until next week, when it will be maneuvered to the Orlan-EVA
7 (from DC-1) viewing position.  [Several failure messages
occurred during the operations, all due to either a known problem
and/or understood and explained.  For example, as temperature of
the MBS mast camera decreases, it eventually hits a point where it
flips to a high value and annunciates an out-of-limits alarm.
 This robotics advisory will be inhibited in the future.
 Another alarm message is annunciated whenever an RWS (robotics
workstation) is powered up and will only return to norm if the RWS is
taken to Active.  No action is required for this or any other
such messages.  One "Command Sequence Error" message
is currently in alarm and will remain in that state until the RWS is
powered up and taken to Active.]

After wake-up and before breakfast, all crewmembers performed the
Med Ops PHS (periodic health status) examination.  As part of
the exam, MO-9 (urinalysis) and blood analysis using the PCBA
(portable clinical blood analyzer) equipment were performed, with the
crew free to do these activities in whatever order they chose, as
long as they were completed before crew breakfast and before the
first exercise session.  After the exam, Treschev stowed the
Urolux equipment. [FE-1 Peggy Whitson acted as primary CMO (crew
medical officer) and CDR Valery Korzun as the backup.  The
ground suggested for Peggy to perform PCBA on Valery and then on FE-2
Sergei Treschev, and Valery to perform PCBA on Peggy last.
 After each test was completed, they could enter the data
directly in the IFEP software on the MEC or wait to enter all three
data sets after everyone had completed PCBA.]

FE-2 Treschev conducted another
Plasma-Crystal-3 experiment run today, first activating the
turbopump, then–if the payload software loaded successfully–running
the plasma experiment in three parts: (1) study of plasma
ãheart-beatä instability at various discharge energy levels; (2)
study of residual charge attached to small particles after
termination of discharge (relaxation); and (3) study of residual
charge attached to large particles after termination of discharge
(relaxation).  In case the software did not load properly,
Sergei was to run two segments of 35 minutes each, with predefined
parameters set at the beginning of each segment and remain constant
during each run.  After experiment completion, he was to test
and calibrate the payload, copy the data to the computer and remove
the videocassette used to record the oscillating plasma cloud.
 Before dinner, the turbopump was to be turned off again.

Yesterday’s test activation of the Elektron oxygen generator after
the Fluids Unit R&R (removal and replacement) performed by CDR
Valery Korzun failed.  Today, Korzun continued its
troubleshooting.  [After first preparing the unit for
subsystem removal and replacement, he again detached it from the
BITS2-12 onboard telemetry system by demating its TMI-connector.
 Assisted by FE-2 Treschev, Valery then removed the BZh Fluids
Unit #3 installed yesterday, and replaced it with the previous Fluids
Unit #4, tagging up with ground specialists via S-band.  The
troubleshooting activity is to be continued tomorrow.]

At 12:01 pm EDT today, MCC-Moscow plans a restart of the TVM
(terminal computer) system in the SM, which has only one of its three
lanes operating.  For this purpose, ISS flight attitude control
will be handed over to the RS at 11:50 am, which will then maneuver
to the restart attitude.  Later, at 1:20 pm, control authority
will be returned to the USOS CMGs for momentum management in XPOP.

The SM TsVM (Central computer) system continues running on all three
subsets in the redundant set.

Today’s SUBSA (solidification using a baffle in sealed ampoules) activities by Peggy Whitson were the completion of
the third sample run (SUBSA-02), sample 2 removal, micro drive
assembly exchange, video tape removal, and MSG (microgravity science
glovebox) powerdown.  Peggy was thanked for all the hard
work she did with the video camera and getting the ground set up for
the Eurotherm controller verification run yesterday.

TVIS  treadmill troubleshooting:   MCC-Houston is
making progress on several fronts.  There are two solid fixes in
work that have a very good chance of making the next Progress launch.
Specialists are working hard to fly a new and improved chassis on 9A
and there is a good chance that that will pan out.  Engineers
are cautiously optimistic that TVIS can safely be used prior to these
fixes arriving on orbit by constraining the TVIS operations to
passive mode with reduced loadings.  Further procedures are
being developed and checked out on the ground unit.

All crewmembers performed their daily physical exercise program, and
Peggy Whitson later transferred the TVIS/RED/CEVIS data, accumulated
in PCMCIA (Personal Computer Memory Card International Association)
cards, to the medical equipment computer for subsequent downlink to

Later today, MCC-H will uplink an upgrade of the onboard MPV (manual
procedures viewer) structure, to facilitate crew access to its stored
voluminous procedures books.

At 4:25 am EDT this morning, Peggy Whitson was scheduled to conduct
an amateur (ham) radio exchange with students at the Kansai Ham
Festival 2002 in Hirakata, Japan.

Daily routine servicing tasks were performed by CDR Korzun (SOSh life
support system maintenance, IMS delta file preparation) and FE-1
Whitson (Lab payloads status checkup).

Solar Beta angle of the ISS orbit is currently 53 deg and

SpaceRef staff editor.